1. (15pts) Events A, B and C are such that P(A) = 0.7, P(B) = 0.6, P(C) = 0.5, P(AnB) = 0.4 , P(AnC) = 0.3, P(BnC) = 0.2, P(AnBnC) Find (a) either B or C happens (b) at least one of A, B, C happens; c) exactly one of A, B, or C happens. 0.1.
Answer #1

0.1 0.3 0.1 0.1 0.1 0.2 0.0 0.1

a) 0.3+0.1+0.1+0.1+0.2+0.1=0.9

b)0.1+0.3+0.1+0.1+0.1+0.2+0.1=1.0

c)0.1+0.1+0.1=0.3